### v1

# Ensemble Watermarks for Large Language Models

Abstract
The rapid advancement of large language models (LLMs) has made it increasingly difficult to distinguish between text written by humans and machines. While watermarks already exist for LLMs, they often lack flexibility, and struggle with attacks such as paraphrasing. To address these issues, we propose a multi-feature method for generating watermarks that combines multiple distinct watermark features into an ensemble watermark. Concretely, we combine acrostica and sensorimotor norms with the established red-green watermark to achieve a 98% detection rate. After a paraphrasing attack the performance remains high with 95% detection rate. The red-green feature alone as baseline achieves a detection rate of 49%. The evaluation of all feature combinations reveals that the ensemble of all three consistently has the highest detection rate across several LLMs and watermark strength settings. Due to the flexibility of combining features in the ensemble, various requirements and trade-offs can be addressed. Additionally, for all ensemble configurations the same detection function can be used without adaptations. This method is particularly of interest to facilitate accountability and prevent societal harm.
Keywords Watermark  $\cdot$ Stylometry  $\cdot$ Ensemble  $\cdot$ Accountability  $\cdot$ LLM
1 Introduction
Large language models (LLMs) made major improvements with the inception of encoder and decoder models using attention in the form of transformers (Vaswani et al., 2017). Further refinements were made to LLMs using model pre-training (Devlin et al., 2019) and very large data sets (Radford et al., 2019). Through these developments, it has continuously become harder for humans to differentiate texts written by LLMs from those written by humans, with machine-generated texts sometimes even fooling humans more often than human-written texts (Zellers et al., 2019). While machine detection techniques exist, they are slowly being outgrown by the progress of LLMs, for example, several detection methods for GPT-2 are already shown to struggle with GPT-3 (Fagni et al., 2021). This trend already causes worries in many aspects that will only increase with the trend of improvement and consistent scaling of LLMs (Kaplan et al., 2020). Several potential cases of misuse already exist for these advanced models (Ray, 2023).
Watermarks attempt to solve this by embedding a secret code into LLM output by modifying logits of the generated tokens during the generation process. However, as we will show later, a watermark with only a single feature has limited resilience against attacks like paraphrasing. To help against this weakness, we introduce an ensemble watermark that combines stylometric watermark features like acrostica and sensorimotor norms with the established red-green watermark feature introduced by Kirchenbauer et al. (2023). Our method is flexible and allows for a diverse set of features, and we draw inspiration from features used in the context of stylometry.
Stylometry is the study of assessing a person’s writing style, with many stylometric features being proposed in literature Neal et al. (2017). These features are like a writer’s fingerprint, including syntax, the vocabulary used, the structure and size of sentences, and other idiosyncrasies of the author. All these features can be detected statistically and used in tasks like authorship attribution (Stamatatos, 2009a). However, classic authorship attribution has been shown to struggle with smaller LLMs, such as GPT-2 Uchendu et al. (2020).
Sensorimotor norms are categories based on human cognition. Perceptual modalities like "hearing" or action effectors like "hand" are well researched in psychology and cognitive semantics Lynott et al. (2020), but there is yet little research in computer science. In our approach, a secret key derived from the generated output selects the sensorimotor category and thus influences the generated words, e.g., for the olfactory category "smells funny" would be preferred over "looks funny".
An acrostic is a text in which the first letter of each sentence can be combined to spell out a hidden message or word. Historically, acrostica have been used by authors by encoding their authorship Johnson (2006) with variations of their names being the hidden acrostic. A famous example of an acrostic is reported in Appendix 3. In our approach, the secret key controls the letters to be used as the first letter of the first word of a generated sentence. The main contributions presented in this paper are:
  • We propose an ensemble watermark approach for LLMs based on changing token logits on a token and sentence-based level to embed novel stylometric features combined with an established red-green watermark.
  • We show that this provides more resilience against paraphrasing attacks for three LLMs and three different parameter settings and provides the best detection rate.
  • We propose a detection method that works for any combination of our ensemble watermark features, even in isolation, without any changes to the function.
The flexible nature of our ensemble watermark allows it to be adapted to different requirements while using the same detection method.
2 Background & Related Work
Table 1: Overview of common stylometric features as identified in literature from Niess and Kern (2024). Their suitability for four distinct ways to integrate watermarks into LLMs are reported based on a self-assessment. A black circle shows the best compatibility, followed by the semi-filled circle. For our ensemble we decided on sensorimotoric words and acrostics next to the established red-green feature, and on logit manipulation for watermarking.
Feature Type Logits Fine-Tuning Prompt Post-Processing
N-Grams
Character Frequency
Vocabulary Richness
Word Distributions
Word Length
Sentence Length
Parts of Speech
Punctuation Frequency
Sentence Complexity
Synonyms
Sensorimotoric Words
Acrostics
Red-Green
Several notable attempts exist to detect machine generated text, albeit without incorporating language and stylometric features. Depending on their functional approach, these efforts can generally be categorized into two primary groups.
2.1 Watermarking Approaches
Post-hoc detection.
In this case, it is assumed that the LLM has already completed text generation sometime in the past. The methods in this group attempt to detect the generated text without employing any watermark or other modifications to either the LLM or the final output. Notable examples of this group are the classifier by OpenAI Kirchner et al. (2023), GPTZero Tian (2023) or DetectGPT (Mitchell et al., 2023) who uses the probability curvature of text sampled from a LLM. DetectGPT utilizes the property that text tends to occupy negative curvature regions of the log probability of the model. This implies that any minor modifications to a sentence will result in a reduction in its log likelihood. This is because an LLM continuously strives to attain the optimal probability for each sentence. More post-hoc methods can be found in the survey by Jawahar et al. (2020). Post-hoc detection methods have the benefit that they can be used on any suspected texts without any prerequisites, but they are susceptible to break by user attacks and become less effective with complex language models Chakraborty et al. (2023)
Watermarking and Red-Green Watermarks.
Watermarking is the technique of hiding information in data that is difficult for others to remove but can be detected by an algorithm to read the hidden information. Some successful watermark implementations already exist, such as Kirchenbauer et al. (2023), Christ et al. (2023), and there is ongoing research by Aaronson (2022, 2023). Specifically, Kirchenbauer et al. (2023), which we also include in our ensemble watermark and all our comparisons, separates tokens into a green and red list, where tokens in the green list get a weight boost to increase their representation in the output. This leads to generated text mostly consisting of words from the green list, while human written text naturally also uses words from the red list. Xiang et al. (2024) address the issue of replacing sensitive words and make use of Sentence-BERT (Reimers and Gurevych, 2019). Zhao et al. (2023) also use red-green watermarking, but instead of dynamically generating the red-green list, use a fixed list to simplify the grouping strategy. Kuditipudi et al. (2024) then test several sampling schemes to improve watermark performance. The Duwak approach Zhu et al. (2024) combines the red-green watermark feature with a sampling scheme to improve watermarking; it does, however, not combine multiple logit watermark features as we do.
2.2 Stylometric Features
There is a broad range of stylometric features previously proposed in literature (Lagutina et al., 2019; Stamatatos, 2009b). They can broadly be categorized into five levels: lexical, syntactic, semantic, structural, and domain-specific features. In Table 1 we collected in previous work Niess and Kern (2024) popular stylometric features reported in literature, as well as additional watermark features that are considered promising.
Sensorimotor Norms.
Winter (2019) define sensory linguistics as the study of how senses and languages are related. The Lancaster Sensorimotor Norms (Lynott et al., 2020) describes a set of 40,000 sensorimotor words together with a crowed-based assessment into 11 dimensions, forming the base of our sensorimotor features. Words can be represented by six perceptual groups (touch, hearing, smell, taste, vision, interoception) and five actions (mouth/throat, hand/arm, foot/leg, head excluding mouth/throat, and torso). Khalid and Srinivasan (2022) use the Lancaster Norms and find that sensorial language might be intentionally used and cannot be considered a random phenomenon. Perceptional features have also been proposed in literature for authorship attribution, specifically in a cross-language setting (Bogdanova and Lazaridou, 2014).
Acrostica.
Stein et al. (2014) considered the task of generating a paraphrased version of an existing text, where the generated text should contain an acrostic. Thereby, they formulated this as a search problem. Shen et al. (2019) utilized a sequence-to-sequence network to generate text incorporating acrostica for English and Chinese. More recently, steganography has been considered to embed secret messages in text using BERT (Yi et al., 2022).
2.3 Watermarking Implementations
LLMs can utilize watermarks, including stylometric ones, in several ways. We present four different approaches in the following. They have different strengths and weaknesses. We provide an overview and estimate how effective each approach is expected to be for different stylometric features in Table 1.
Logits Manipulation.
In our approach, we have opted for directly manipulating the logits of the LLM tokens generated. Logits are the raw output values of a machine learning model before applying an activation function like softmax. They represent the unnormalized score or prediction for each token. This manipulation gives considerable control over how often single features are generated, though the difficulty lies in finding the correct logits that produce the desired features. An example of a watermark of this type is the work by Kirchenbauer et al. (2023).
Fine-tuning.
A standard solution is to conduct additional training to change the output of an LLM. This, of course, also works with watermarks, though it is less controlled than other methods. For example, the written material of an author with a known writing style could be used to fine-tune an LLM to produce text similar to that author Li et al. (2023). The problem is that it is difficult to control what is learned and what is not. The model might learn not only the desired watermark features but also overfit to other features, which can lead to a loss of generalization, domain mismatch, and, in general, reduced robustness for outputting text.
Prompt Engineering.
Another option is to use the innate capability of LLMs and to engineer a system prompt that tells the model what features are desired Zhou et al. (2023). For example, it is possible to tell contemporary LLMs to use only certain letters in their sentences OpenAI ChatGPT (2024). The challenges here are creating the right prompt, defending against users writing their own prompt, and determining what prompts the language model can comprehend and adhere to. In general, a more powerful model is more suitable for this method.
Post-Processing.
The last possibility is to post process the text after generation is finished. This is how early attempts were made by Topkara et al. (2005) and Atallah et al. (2001) when LLMs were not yet developed. This was popular for embedding watermarks at the time, for example, for copyright or document integrity. Although the generative power of LLMs has made this less attractive, it can still be useful for some simple features like synonym replacement that do not require the capability to generate new sentences.
3 Method
Our approach consists of two main parts: 1) the watermark generation process based on manipulation of logits controlled via dynamic keys, and 2) a test procedure for detecting an existing watermark based on statistical tests.
3.1 Watermark Generation
The generation algorithm modifies the logits of the language model during text generation to embed the features forming the watermark. It adjusts token probabilities to make certain words more likely, based on: The acrostic pattern, by boosting tokens that start with specific letters. Sensorimotor words, by boosting tokens associated with a target sensorimotor class. The red-green mechanism, by adjusting token probabilities based on a dynamically generated green list. We follow the procedure proposed in the original paper Kirchenbauer et al. (2023) for splitting the red-green list. For the other features, a secret key is required, as described below.
Algorithm 1 Watermarked Text Generation
[…]
2:  Set $\delta_{\text{acro}},\delta_{\text{senso}},\delta_{\text{redgreen}}$
3:  while not done generating do
4:     Get current logits from the model
5:     if starting a new sentence then
6:        Adjust logits for acrostic boosting
7:     else
8:        Adjust logits for sensorimotor boosting
9:     end if
10:     Generate green list based on last token
11:     Adjust logits for red-green mechanism
12:     Sample next token from adjusted logits
13:     Update the secret key based on last word or sentence
14:  end while
Secret Key Generation.
A secret key is maintained throughout the generation process to control most of the features, i.e., it determines the sensorimotor class and the letter for the acrostic pattern. The key is updated based on the last word and the last sentence using secure hash functions. Both words and sentences are hashed using the same base function. Given a word $w$, the hash function maps it to an integer directly within a specified range $[a,b]$. Given a sentence $s$, we first lemmatize and remove stopwords and punctuation to get a sentence $s^{\prime}$. The hash function is applied to $s^{\prime}$ to generate an integer within a range $[a^{\prime},b^{\prime}]$.
[…]
$\left.\mod(b-a+1)\right)+a$
Logits Adjustment.
During generation, the logits (raw scores before softmax) are adjusted to boost the probability of specific tokens. For example, if a new sentence is started, the initial token is boosted according to the target acrostic letter. For each token $t$:
[…]
Otherwise, boost tokens that are associated with the current sensorimotor class. For each token $t$:
[…]
The red-green mechanism is based on the work by Kirchenbauer et al. (2023). A green list is generated based on the last token $t_{\text{last}}$. A random number is then seeded by a generator with: $\text{seed}=\text{hash}(t_{\text{last}})$. This is used to generate a random permutation of the vocabulary, where the first $\gamma\cdot V$ tokens are selected as the green list, where $V$ is the vocabulary size and $\gamma$ is a predefined proportion (e.g., $0.5$) For tokens in the green list: $\text{logits}[t]+=\delta_{\text{redgreen}}$
It is important to ensure consistent tokenization between generation and detection and appropriately handle special tokens (e.g., BOS, EOS). Streaming generation is used to update the secret key and adjust logits at each step during generation. To detect sentence boundaries, punctuations are used (e.g., ., !, ?).
We chose a relatively large weight boost for acrostica since the beginning of the sentence is more flexible, and the rest of the sentence can easily adapt to this change. We chose a small weight for sensorimotor words so that the model would not be overly biased.
All features have in common that their strength is controlled via a $\delta$ parameter. In the evaluation we study the impact of these parameters on the generated text.
While we opted for fixed values for the weights, they could also be chosen dynamically, depending on how long it has been since a desired feature was chosen or on the distribution of the current weights.
3.2 Watermark Detection
The detection of the watermark works similarly to that of the generation. A secret key is maintained the same way, but instead of modifying logits, the generated token is compared based on the key, and the probability of that token occurring is calculated.
Algorithm 2 Watermark Detection Algorithm
1:  Input: Text $T$
2:  Initialize: Load sensorimotor norms and class frequencies from the corpus
[…]
4:  Initialize counters for acrostic matches $k$, total checks $n$
5:  Initialize counters for sensorimotor matches per class $k_{c}$, total words per class $n_{c}$
6:  Initialize variables for red-green detection: total transitions $T$, green tokens $G$
7:  Split text $T$ into sentences $S=[s_{1},s_{2},\dots,s_{N}]$
8:  for each sentence $s_{i}$ in $S$ do
9:     Split $s_{i}$ into words $W=[w_{1},w_{2},\dots,w_{m}]$
10:     if $i>1$ then
[…]
12:        Compare with first letter of $w_{1}$
13:        Increment $k$ if match found
14:     end if
15:     for each word $w_{j}$ in $W$ do
[…]
18:        if $w_{j}$ is in sensorimotor dictionary then
19:           $n_{c}\leftarrow n_{c}+1$
20:           if $w_{j}$ belongs to class $c$ then
21:              $k_{c}\leftarrow k_{c}+1$
22:           end if
23:        end if
24:        Red-Green Detection: Update green list based on previous token
25:        Increment $G$ if $w_{j}$ is in green list
26:     end for
27:  end for
28:  Calculate Probabilities: Compute acrostic probability $P_{\text{acrostic}}$
29:  Compute sensorimotor probability $P_{\text{sensorimotor}}=\prod_{c}P_{\text{sensorimotor},c}$
30:  Compute red-green probability $P_{\text{redgreen}}$
31:  Compute Final Score:
[…]
32:  Output: final_score
Figure 1: Plot of the Z-Scores of different configurations of the watermark ensemble (Llama 3.1 8B, medium strength watermark). The human and red-green results serve as baselines. Each watermark configuration has 400 unique samples. The combination of all features achieves the best detection rate of 97.75%. The combination of acrostic and red-green might be attractive for its more moderate increase in perplexity (+0.85), while still achieving a detection rate of 95.75%. The red-green baseline achieves a detection rate of 83.00%.
The probability of detecting an acrostic watermark is calculated using the formula:
$P_{\text{acrostic}}=1-\sum_{i=0}^{k-1}\binom{n}{i}p^{i}(1-p)^{n-i}$
where $n$ is the number of acrostic checks (total sentences minus one), $k$ is the number of correct matches observed, and $p=\frac{1}{26}$ represents the probability of a random match.
The probability of detecting a sensorimotor watermark can use class-specific frequencies from the Google N-gram corpus or evenly split probabilities. For each class $c$, the baseline probability $p_{c}$ is calculated. Given $k_{c}$ matches out of $n_{c}$ words for class $c$, the probability is:
$P_{\text{sensorimotor},c}=1-\sum_{i=0}^{k_{c}-1}\binom{n_{c}}{i}p_{c}^{i}(1-p_{c})^{n_{c}-i}$
The combined probability across all classes is:
$P_{\text{sensorimotor}}=\prod_{c}P_{\text{sensorimotor},c}$
The probability calculation for the red-green watermark involves the following parameters: $T$, which represents the total number of transitions; $G$, the number of tokens in the green list observed; and $\gamma=0.5$, the probability parameter. The expected number $E$ and variance Var are:
$E=\gamma T,\quad\text{Var}=T\gamma(1-\gamma)$
The Z-score and probability is calculated as:
$Z=\frac{G-E}{\sqrt{\text{Var}}},\quad P_{\text{redgreen}}=1-\Phi(Z)$
where $\Phi$ is the cumulative distribution function of the standard normal distribution. Assuming independence, the final score is:
[…]
4 Evaluation
In this section, we describe the experiments conducted, assess the resilience against a paraphrasing attack, and conduct an ablation study to analyze the impact of text length.
4.1 Implementation Details.
We used three LLMs, Llama 3.1 8B (Dubey et al., 2024), Llama 3.2 3B, and Mistral 7B Jiang et al. (2023). We utilized the logits processor of Hugging Face Wolf et al. (2020) to manipulate the generation. To create prompts with human baselines, we randomly select texts from the C4 RealNewsLike dataset (Raffel et al., 2019), trim a fixed length of tokens as "baseline" completions from the end, and treat the remaining tokens as the prompt.
We create a weak, middle, and strong parameter setting to test different parameters for the features. The medium setting [$\delta_{senso}=2.5,\delta_{acro}=20.0,\delta_{redgreen}=2.0$] has 400 samples per configuration, while the strong [$\delta_{senso}=5.0,\delta_{acro}=40.0,\delta_{redgreen}=10.0$] and weak [$\delta_{senso}=1.0,\delta_{acro}=10.0,\delta_{redgreen}=1.0$] setting contain 300 samples per configuration.
We use a level of $\alpha=0.05$ to determine the statistical significance that a watermark can be recovered. This allows us to calculate a detection rate showing how many samples would have been detected as a watermark with this $\alpha$ threshold.
Figure 2: Llama 3.1 8B tested with three different parameter settings. Higher parameter values for the $\delta$ weights lead to lower Z-Scores (improved detectability) and an almost linear increase in the perplexity. The acrostic feature stands out for having the least impact on perplexity, with the ensemble of all features providing the best detectability for each parameter setting.
4.2 Results
Figure 1 provides an overview of combinations of features for Llama 3.1 8B, with a lower Z-Score indicating improved detectability. Overall, combinations of watermark features have a higher detection rate than single features. Of the single features, the red-green feature provides the best detectability.
The combination of all features has the best detection rate, with an increase of 17.78% compared to the red-green feature. We also tested for statistical significance in Figure 3. This finding is consistent across all experiments we conducted and reported in Appendix A for all settings.
Figure 3: Heatmap of adjusted p-values from pairwise Mann-Whitney U tests between configurations (Bonferroni correction applied) for Llama 3.1 8B with the medium parameter setting. Lower adjusted p-values indicate a significant difference after correction for multiple comparisons.
Configuration Human Senso Acro Red-Green Senso + Acro Senso + Red-Green Acro + Red-Green All Three
Llama 3.1 8B (Strong) 0.34 80.41 28.52 49.14 89.35 84.19 55.67 95.19
Llama 3.2 3B (Strong) 0.97 85.11 31.39 54.05 90.61 90.29 64.08 95.79
Llama 3.1 8B (Medium) 2.42 58.06 28.23 34.14 73.92 70.70 46.77 82.53
Mistral 7B (Medium) 1.44 54.87 42.60 23.47 69.31 58.84 40.07 73.65
Llama 3.2 3B (Weak) 1.32 26.07 27.72 29.04 38.61 43.56 35.97 44.88
Table 2: Detection rate of different LLM configurations after a paraphrasing attack changing at least 10% of each text. All three watermark features combined are consistently the best configuration. Interestingly, Mistral 7B has better results for the acrostic feature since it tended to produce more numerous shorter sentences.
The scatter plot in Figure 2 additionally shows the three considered parameter settings (weak, medium, strong), together with the increase in perplexity. Higher perplexity reflects more substantial deviations in the output distribution. For all parameter settings, combining all features provides the best detectability but with an associated increase in perplexity. Interestingly, the acrostic feature is associated with the smallest increase in perplexity, i.e., the most minor influence on the distribution of overall tokens.
Experiments on Paraphrasing Attacks.
We conducted experiments to assess the vulnerability of the various watermarking configurations via paraphrasing attacks. This experiment involves tokenizing the watermarked text, then iteratively replacing one word with a <mask>, and using T5 (Raffel et al., 2019) to generate candidate replacement sequences via beam search. If one candidate differs from the original string, the attack succeeds, replacing the span with the new text. This is repeated until at least 10% of the watermarked text is replaced. The results can be found in Table 2 and show that the combination of all three features has the highest detection rate after the paraphrasing attack for all of the considered LLMs and parameter settings. Llama 3.1 3B with the strong parameter setting retains over 95% detection rate, in contrast to the red-green feature in isolation, which drops to 49%. Here, the sensorimotor feature stands out, being more resilient to this type of attack retaining over 80% detection rate. The acrostic feature alone does not appear to be resilient enough to the attack and only achieves a low detection rate. However, it still contributes to overall watermark resilience, which can be seen in the improvement of the ensemble of all features against using only the combination of sensorimotor and red-green feature. Overall, using multiple features together contributes to much greater resilience against paraphrasing attacks.
4.3 Ablation Study Text Length
Figure 4: For the ablation study, we study the influence of the number of output sentences on the detection rate. We selected Llama 3.1 8B as language model, and present results for the weak, middle and strong parameter settings. The ensemble of all three features provides the best overall detection rate with the fewest sentences.
To further analyze the influence of text length on the contribution of each feature, we do incremental sentence partitioning, where each version represents a progressively shorter subset of each text, starting with the total sentence count and reducing by one until reaching a single sentence. In Figure 4, the detection rate of each feature can be seen for three different watermark configurations. While acrostics are not as sensitive to differences in the $\delta$s, both sensorimotor and red-green features are associated with a jump in detection rate once an increase in the weights is applied to the logits. It can also be seen that a combination of all three features reaches a high detection rate earlier than other configurations and that the weak setting even beats some of the medium-strength configurations, hinting that for specific requirements and trade-offs specific combinations in the ensemble are best suited.
5 Conclusion
We presented a novel ensemble watermark method for large language models, particularly generative transformer models. Out of the possible approaches for integrating watermarks in LLMs we opted for manipulating the probability directly when generating the tokens. Following existing work, we generate a key dynamically derived directly from the generated text to control the watermark features. Out of many possible stylometric features, we focused on two features for our evaluation, the acrostic and sensorimotor norms, and combined them with an established method from literature, red-green lists.
For the experiment, we selected three different LLMs and three levels of watermark strength. In the evaluation, we found that each of the three features has different characteristics in relation to the detectability and perplexity of the generated sentences.
The acrostic feature has the most negligible impact on perplexity, but its performance depends on sentence number and length and has less resilience against paraphrasing than other features. The sensorimotor feature is similar in watermark performance and perplexity impact as the established red-green feature but shows increased resilience against paraphrasing. On the other hand, the red-green feature shows a balanced performance and combines well with the other (stylometric) features.
Overall, we propose a flexible and resilient ensemble watermark for text that works with short text lengths and does not require a different sampling strategy, expensive additional model training, or an LLM for testing. Since our method allows for many types of key generation and (stylometric) features, an exploration of further combinations, including watermark sampling strategies, will be part of the future work.
Appendix A Appendix
Table 3: Example of an acrostic written by Arnold Schwarzenegger in a letter to the members of the California State Assembly, which has been used in literature for illustration before [Stein et al., 2014].
To the Members of the California State Assembly:
I am returning Assembly Bill 1176 without my signature.
F or some time now I have lamented the fact that major issues are overlooked while many
u nnecessary bills come to me for consideration. Water reform, prison reform, and health
c are are major issues my Administration has brought to the table, but the Legislature just
k icks the can down the alley.
Y et another legislative year has come and gone without the major reforms Californians
o verwhelmingly deserve. In light of this, and after careful consideration, I believe it is
u nnecessary to sign this measure at this time.
Sincerely,
Arnold Schwarzenegger